q varies jointly as m and n.if q=24 when m =4 and n =3, find m when q=36 and n=4
Answer by Alan3354(69443)

q varies jointly as m and n.if q=24 when m =4 and n =3, find m when q=36 and n=4
-----------
q = k*m*n where k is a constant.
24 = k*4*3
k = 2
------------
36 = 2*m*4
m = 4.5